A hotel has 150 rooms. A room with extras costs $80 a night, and a room w/o costs $60 a night. On a night that the hotel is full, the total revenue is $10,000. How many of each room does the hotel have?
Found 2 solutions by ewatrrr, richwmiller:
Answer by ewatrrr(24785)

x the number of $80 rooms
80x + 60(150-x) = 10,000
x = (10,000 - 9,000) / 20 = 50, $80 rooms and 100 $60 rooms
And..checking
$4000 + $6000 = $10,000

Answer by richwmiller(17219)

x+y=150,
80*x+60*y=10000
x=150-y
80*(150-y)+60*y=10000
12000-80y+60*y=10000
-20*y=-2000
y=100
x=150-y
x=50 rooms at $80 ($4000) and y=100 rooms at $60 ($6000)
check
50+100=150
80*x+60*y=10000
80*50+60*100=10000
4000+6000=10000
10000=10000
ok
codetick